There are 10 multiple choice questions on a test , with 5 possible answers to each question. What is probability of getting 6 out of 10 correct ?
1
Expert's answer
2020-03-13T12:16:22-0400

Total number of ways of answering the 10 questions is 510

Number of ways of answering 6 out of 10 questions correct is "{10\choose 6}\cdot4^4" = "210\times 256"


Probability of getting 6 out of 10 correct questions is "\frac{210\times256}{5^{10}} = \frac{10752}{5^{9}} = 0.0055"